1. The momentum of a garbage truck that is 1.20 × 104 kg and is moving at 10.0 m/s is
  2. At what speed would an 8.00-kg trashcan have the same momentum as the truck?
1
Expert's answer
2021-03-04T17:24:01-0500

1)

"p=mv=1.2\cdot10^4\ kg\cdot10\ \dfrac{m}{s}=1.2\cdot10^5\ \dfrac{kgm}{s}."

2)

"v=\dfrac{p}{m}=\dfrac{1.2\cdot10^5\ \dfrac{kgm}{s}}{8\ kg}=1.5\cdot10^4\ \dfrac{m}{s}."